A bicyclist races at 20 meters per second and beats her record by 12 seconds. The bicyclist races again at 18 meters per second and is over her record by 10 seconds. What was the distance of the race?
Answer by lwsshak3(11628)

A bicyclist races at 20 meters per second and beats her record by 12 seconds. The bicyclist races again at 18 meters per second and is over her record by 10 seconds. What was the distance of the race?
***
let x=distance of race
let t=record travel time
travel time=distance/speed
x/18=t+12
x/20=t-10
x/18-x/20=t+12-t+10=22
x/18-x/20=22
20x-18x=18*20*22
2x=792
x=396
distance of race=396 m